# Robust Ladder Climbing with a Quadrupedal Robot

### v1

@@ III-A Simulation Results
Figure 4: A simulated robot was assessed for its ability to traverse ladders of different configurations. Shown is a composite image generated in simulation of the longest ladder tested: $L_{len}=\textnormal{3~{}m}$ at $L_{\theta}=\textnormal{90}^{\circ}$.
@@
The student policy was evaluated in simulation across various $L_{\theta}$, $L_{radius}$, and in the presence of noise and external disturbances. Noise was added to provide a more realistic evaluation of the robustness of the policy. During the evaluation, $L_{len}$ was randomly sampled between 1-3 m, $L_{width}$ between 1.0-1.25 m, and $L_{space}$ between 27.5-32.5 cm. We sampled 50 different ladder configurations and averaged the results across 3072 agents simulated for 15 s. To more fairly compare the hook end-effector to the traditional ball-foot, the ladder rungs were made cylindrical during evaluation, rather than the elliptic cylinders used during training. An example evaluation is shown in Fig. 4.
@@
The robot with the hooked end-effectors achieved an average 96% success rate across the configurations tested (see Fig. 5 for the success rate for each configuration). The robot successfully climbed ladders in the presence of added noise and external disturbances, such as random pushes applied to the base every 5 s, sampled from a normal distribution with a standard deviation of 1 m/s. In the case of $L_{radius}=\mathrm{3.5}~{}\textnormal{cm}$, which is larger than the opening radius of the hook, we observed a steady drop in performance at steeper $L_{\theta}$ because the robot could not stabilize itself with the forelegs. As $L_{\theta}$ increases, the robot’s applied forces switch from mostly compressive to a mix of compressive and tensile forces needed to stabilize the center of mass.
@@
We compared the policy trained with the hook end-effector against another policy trained using the traditional ball-foot design, finding a clear performance gap across all configurations and a notable performance drop at steeper $L_{\theta}$. In fact, the average success rate across all configurations was only 81% for the ball end-effector (Fig. 5). Interestingly, the performance of the ball end-effector improved with decreasing $L_{radius}$ across all $L_{\theta}$ tested in simulation. We observed that the policy can exploit the small crease between the ball foot and the shank to weakly anchor the robot on the rungs at these smaller values of $L_{radius}$. Moreover, with decreasing $L_{radius}$, it became easier for the robot to maneuver its knees between the rungs, a movement strategy the policy relies on heavily in the absence of hooks.
@@
To better understand the relative benefits of the hook end-effector, we ran noise-free simulations over the configurations in Fig. 5. The ball and hook end-effectors both yielded success rates greater than 99% in all configurations. Thus, the hook does not enable climbing in simulation per se, but furnishes stability to generalize over uncertainty and enables smooth and rapid climbing behavior. We see this as a critical factor for robust real-world performance.
@@
Figure 5: Climbing success rate in simulation across various ladder inclinations and rung radii, juxtaposing performance of the ball and hooked end-effector. The agents were evaluated in the presence of noise and external disturbances.
III-B Real World Results
@@
The policy was deployed zero-shot on an ANYmal D robot (ANYbotics AG) without further fine-tuning. A motion capture system was used to estimate the ladder pose, $p_{l}$, along with $L_{\theta}$. Other elements of the ladder state, $s_{l}$, were measured directly and input as scalars to the policy. In our setup, the reference ladder had parameters $L_{width}=\mathrm{1}~{}\textnormal{m}$, $L_{len}=\mathrm{1.8}~{}\textnormal{m}$, $L_{space}=\mathrm{30}~{}\textnormal{cm}$ (with five cylindrical rungs), and $L_{radius}=\mathrm{2.5}~{}\textnormal{cm}$. The ladder was placed at various $L_{\theta}$ against a stack of wooden boxes.
@@
Figure 6: Representative sim2real experiment. Key snapshots highlighted: ladder mounting, midway point, approaching the dismount, and dismount. Plots show robot base position and orientation data over time while ascending a ladder with $L_{\theta}=\textnormal{80}^{\circ}$. Consistency between simulation and reality is furthermore observed throughout the experiment through low deviation between Real and Sim curves.
Performance at varying $L_{\theta}$: At $L_{\theta}$ values of 70 and 80, the policy succeeded in four of four tests in each configuration. At $L_{\theta}=\textnormal{90}^{\circ}$, the policy succeeded in two of the three tests. During the third test at 90, unmodeled standoffs on the base of the robot were observed to collide with the topmost rung and prevented the robot from climbing up. We decided not to proceed with further tests at 90 due to this obvious sim2real gap.
@@
A representative sim2real experiment with a ladder at $L_{\theta}=\mathrm{80^{\circ}}$ is shown in Fig. 6. Snapshots of key moments highlight the close similarity between simulation and real experiments. Joint positions and environmental contacts during mounting, passing the midway point, and dismounting–culminating in a rapid roll and flick of the back right leg–closely mirror the observed simulation. The position and orientation of the robot base frame between simulation and reality had an average root mean squared error of 11 cm and 0.18 rad, respectively, over the entire trajectory.
@@
Climbing speed: We determined the robot’s average climbing speed from video footage of real experiments. We performed calculations using the time elapsed between the moment the robot first touched the ladder and the moment it fully dismounted. Over ten successfully tested ladder configurations, the average speed was 0.44 m/s (0.51 bl/s) with a standard deviation of 0.16 m/s. For context, our method yields 232$\times$ faster ladder climbing speeds than the existing state-of-the-art quadruped robot in reference [16].
@@
Unmodeled perturbations: The policy was also tested for robustness to unmodeled perturbations. In particular, we tied a rope to the front part of the chassis, back part of the chassis, or front left foot of the robot, and pulled at different moments during its ascent. Fig. 7 depicts an experiment in which the rope was tied to the robot’s front left foot. Analysis of the reaction forces over a characteristic window of time during which the robot was being pulled reveals that the robot can switch from pushing down on the rung (positive $F_{z}$) to generating tensile forces with the hook that it uses to brace itself. In other trials, the policy was found to exhibit persistent recovery and retry behaviors while being pulled from points on the base (Supplementary Video).
@@
Figure 7: The robot successfully climbs ladders even in the presence of unmodeled perturbations. Shown is an excerpt from a test at 80 with a rope tied to the robot’s front left foot. We pulled the rope at various points during the ascent. Due to the shape of the hook, the robot could successful anchor itself, generating tension forces between its body and the rungs with its front right foot (seen as increase of $F_{x}$ and $F_{y}$ to the positive region of the plot and drop of $F_{z}$ to the negative region).
III-C Analogy to human ladder climbing
In humans, a diagonal gait pattern is the most common method of ladder ascent. Upward thrust is generated with the hind legs, and the hands primarily maintain dynamic stability in a direction perpendicular to the ladder uprights [9, 27]. Our experimental results suggest that quadrupedal robots’ ladder ascent motions mirror this pattern, with the knees or hooked portions of hind legs generating upward thrust and the hooked forelegs primarily maintaining stability.
The duality of modes observed between humanoid and quadruped ladder climbing suggests that there exists a unifying description for the functional morphologies and control policies of humanoid and quadrupedal robots, and that they may share a large portion of the same task space. Indeed, though we humans start our lives as quadrupeds, we become primarily bipedal, sometimes reverting to a quadrupedal configuration for certain tasks (like ladder climbing).
IV Conclusion
We demonstrated quadrupedal ladder climbing via a new robot end-effector design and a complementary RL-derived control policy. Our method inherently accomplishes transitions between walking and climbing locomotion modes, eliminating the need to deploy separate policies for separate modes of locomotion. The gaits and climbing behavior exhibited by the policy are completely emergent, yet interestingly, are mechanically similar to human climbing.
@@ IV Conclusion
A comprehensive assessment of performance in simulation revealed an overall 96% success rate at climbing ladders, even with disturbances. Subsequent sim2real experiments validated that the proposed approach elicits robust and reliable policies on robot hardware. Crucially, compared to the traditional ball foot, we found that the hook end-effector furnished the stable shape needed to anchor the robot to the rungs and hang in tension with the center of mass outside of the support triangle, enabling it to traverse steeper ladders and withstand unmodeled perturbations to the base and feet. This finding emphasizes the importance of synergies between geometry and control policy for enhancing robot capabilities.
Future work will focus on realizing quadrupeds that are capable of climbing up and down ladders. Integrating different sensing modalities into the student training pipeline, such as depth camera images, will remove need for the motion capture system and facilitate climbing ladders at industrial sites outside of the lab. Lastly, the present hook design is not optimized, but rather chosen heuristically. There is an exciting frontier of research concerning co-optimization of morphology and control policy to accomplish multiple advanced skills with a single system.
V ACKNOWLEDGMENTS
@@
We would like to thank Takahiro Miki for assisting with demonstrations and brainstorming, and Fabian Tischauser for help with hardware. This work was supported by The Branco Weiss Fellowship - Society in Science, administered by ETH Zurich.

### v2

@@ IV Results
Figure 4: A simulated robot was assessed for its ability to traverse ladders of different configurations. Shown is a composite image generated in simulation of the longest ladder tested: $L_{\mathrm{len}}=\textnormal{3~{}m}$ at $L_{\mathrm{\theta}}=\textnormal{90}^{\circ}$.
@@
Simulation Results: The student policy was evaluated in simulation across various $L_{\mathrm{\theta}}$ and $L_{\mathrm{radius}}$. The policy was evaluated in the presence of observation noise and external disturbances, to provide a more realistic evaluation of policy robustness. For evaluation, $L_{\mathrm{len}}$ was randomly sampled between 1-3 m, $L_{\mathrm{width}}$ between 1.0-1.25 m, and $L_{\mathrm{space}}$ between 27.5-32.5 cm across 50 different ladders. Agents who reached the goal without terminating in under 15 s were marked as successful and the results were averaged among 3072 agents. The rungs were made purely cylindrical during the evaluation, rather than the elliptic cylinders used during training. An example evaluation is shown in Fig. 4.
@@
The robot with the hooked end effector achieved an average success rate 96% across all configurations tested (see Fig. 5). The robot successfully climbed ladders in the presence of added observation noise and external disturbances, such as random velocity offsets applied to the base every 5 s, sampled from a normal distribution with a standard deviation of 1 m/s. In the case of $L_{\mathrm{radius}}=\mathrm{3.5}~{}\textnormal{cm}$, which is larger than the opening radius of the hook, we observed a steady drop in performance at steeper $L_{\mathrm{\theta}}$ as the agent could no longer stabilize itself with the front legs. As $L_{\mathrm{\theta}}$ increases, the robot’s applied forces switch from mostly compressive to a mix of compressive and tensile forces needed to stabilize the center of mass.
@@
We compared the performance of the hook to a policy trained with the traditional ball-foot, observing a clear performance gap across all configurations and a notable performance drop at steeper $L_{\mathrm{\theta}}$. The average success rate across all configurations was only 81% for the ball-foot (Fig. 5). Interestingly, the performance of the ball-foot design improved with decreasing $L_{\mathrm{radius}}$ across all configurations. We observed that the policy can exploit the small crease between the foot and the shank to weakly anchor the robot on the rungs at smaller $L_{\mathrm{radius}}$. Moreover, with decreasing $L_{\mathrm{radius}}$, it became easier for the robot to maneuver its knees between the rungs, a movement strategy the policy relies on heavily in the absence of the hooked end effector.
@@
To better understand the benefits of the hooked end effector, we ran noise-free evaluations over the configurations in Fig. 5. In this case, both designs yielded success rates greater than 99% in all configurations. From this, we conclude that the hook furnishes stability to generalize over uncertainty and enables smooth and rapid climbing behavior—critical factors for robust real-world deployment.
@@
Figure 5: Climbing success rate in simulation across various ladder inclinations and rung radii, juxtaposing performance of the ball and hooked end effector. The agents were evaluated in the presence of noise and external disturbances.
@@
Real World Results: The control policy was deployed zero-shot on an ANYmal D robot (ANYbotics AG) without further fine-tuning. A motion capture system was used to estimate the ladder pose, $p_{l}$, along with $L_{\mathrm{\theta}}$. Other elements of the ladder state, $s_{l}$, were measured directly and input as scalars to the policy. In our setup, the reference ladder had parameters $L_{\mathrm{width}}=\mathrm{1}~{}\textnormal{m}$, $L_{\mathrm{len}}=\mathrm{1.8}~{}\textnormal{m}$, $L_{\mathrm{space}}=\mathrm{30}~{}\textnormal{cm}$ (with five cylindrical rungs), and $L_{\mathrm{radius}}=\mathrm{2.5}~{}\textnormal{cm}$. The ladder was placed at various $L_{\mathrm{\theta}}$ against wooden boxes. At $L_{\mathrm{\theta}}$ values of 70 and 80, the policy was successful in four of four tests in each configuration. At $L_{\mathrm{\theta}}=\textnormal{90}^{\circ}$, the policy was successful in two of the three tests. During the third test, unmodeled standoffs on the underside of the robot were observed to collide with the top rung. We decided not to proceed with further tests at 90 due to this obvious sim2real gap.
@@
Figure 6: Representative sim2real experiment. Key snapshots highlighted: A) ladder mounting, B) midway point, C) approaching the dismount, and D) dismount. Plots show robot base position and orientation data over time while ascending a ladder with $L_{\mathrm{\theta}}=\textnormal{80}^{\circ}$. Consistency between simulation and reality is furthermore observed throughout the experiment through low deviation between Real and Sim curves.
@@
A representative sim2real experiment with a ladder at $L_{\mathrm{\theta}}=\mathrm{80^{\circ}}$ is shown in Fig. 6. Snapshots of key moments during climbing highlight the close similarity between simulation and real experiments. Joint positions and environmental contacts during mounting, passing the midway point, and dismounting—culminating in a rapid roll and flick of the back right leg—closely mirror the observed simulation. In this particular example, the position and orientation of the robot base had an average root mean squared error of 11 cm and 0.18 rad, respectively, over the entire trajectory.
@@
The robot’s average climbing speed was determined from video footage of the real experiments, using the time elapsed from first contact to the moment when the robot fully dismounted from the ladder. Over ten successful runs, the average speed was found to be 0.44 m/s (0.51 bl/s) with a standard deviation of 0.16 m/s. For context, our method yields 232$\times$ faster ladder climbing speeds than the existing state-of-the-art quadruped robot in reference [16].
@@
The policy was also tested for robustness to unmodeled perturbations. A rope was tied to different parts of the robot’s chassis and feet, and pulled at different times during ascent. Fig. 7 shows an experiment in which the rope was tied to the robot’s left front foot. Analysis of the reaction forces over a characteristic window of time during which the robot was being pulled reveals that the robot can switch from pushing down on the rung (positive $F_{z}$) to generating tensile forces with the hook that it uses to brace itself. In other trials, the policy was found to exhibit persistent recovery and retry behaviors while being pulled from points on the base (Supplementary Video).
@@
Figure 7: The robot successfully climbs ladders even in the presence of unmodeled perturbations. Shown is an excerpt from a test at 80 with a rope tied to the robot’s front left foot. We pulled the rope at various points during the ascent. Due to the shape of the hook, the robot could successfully anchor itself, generating tension forces between its body and the rungs with its front right foot (seen as increase of $F_{x}$ and $F_{y}$ to the positive region of the plot and drop of $F_{z}$ to the negative region).
V Conclusion
We demonstrated quadrupedal ladder climbing via a new robot end effector design and a complementary RL-derived control policy. The end effector’s design allows for seamless walking on both nominal and rough terrain, while the controller naturally transitions between walking and climbing without requiring separate locomotion policies. Both the gait patterns and climbing behavior emerge entirely from the learned policy, effectively exploiting the custom hook-shaped end effector.
@@ V Conclusion
Performance assessments in simulation revealed an overall 96% success rate at climbing ladders, even with disturbances. Subsequent sim2real experiments validated that the proposed approach elicits robust and reliable policies on robot hardware. Crucially, compared to the traditional ball foot, the hook end effector furnished the stable shape needed to anchor the robot to the rungs and hang in tension with the center of mass outside of the support triangle, enabling it to traverse steeper ladders and withstand unmodeled perturbations to the base and feet. This finding emphasizes the importance of synergies between geometry and control policy for enhancing robot capabilities. Future work will focus on realizing quadrupeds that are capable of climbing up and down ladders. Adding sensing modalities to the student training pipeline, such as depth camera images, will facilitate climbing ladders at industrial sites outside of the lab that are free standing, contain flat rungs, or have tapered designs.
VI ACKNOWLEDGMENTS
@@
We thank Takahiro Miki and Fabian Tischhauser for their help. This work was supported by The Branco Weiss Fellowship - Society in Science, administered by ETH Zurich.